# LaCATODA 2026: The 10th Linguistic and Cognitive Approaches to Dialog Agents Workshop at the 40th AAAI conference, Janurary 20–27, 2026, Singapore Expo

### v3

@@ 4.5 Limitations and Future Work
Domain adaptability: Our current evaluation focuses on a specific domain of vehicle compliance-related customer service. Future work should systematically investigate the framework’s transferability to broader multi-turn dialogue scenarios, including complex task-oriented dialogues and cross-domain interactions where conversations span multiple service areas.
@@ 5 Conclusion
To address the challenges of poor performance in multi-turn customer service dialogue systems, we propose CID-GraphRAG, a novel framework that integrates intent-driven graph structures with semantic similarity retrieval mechanisms. Our approach leverages knowledge graphs to enhance reasoning capabilities and enable more precise retrieval, while systematically incorporating dialogue history to improve contextual understanding and response coherence, thereby significantly boosting answer quality and relevance.
Experimental results demonstrate that CID-GraphRAG significantly outperforms both semantic-only and intent-only baseline approaches across all evaluation criteria, achieving a $58\%$ relative improvement in response quality over conventional RAG methods according to LLM-as-judge head-to-head evaluations. These findings establish CID-GraphRAG as an effective framework for advancing conversational AI in knowledge-intensive domains where both contextual relevance and goal-oriented progression are essential. In the future, we will explore the adaptability of intent hierarchy, integration of reinforcement learning and extend its applicability to broader multi-turn dialogue scenarios.
Appendix A Intent Recognition Framework and Implementation
@@
Scenario 1. Refusing a proposed appointment time Assistant: I see your vehicle compliance sticker expired last week. We have an opening tomorrow at 2:00 PM to replace it and bring your vehicle back into compliance. User: That doesn’t work for me. Primary intent ($I^{\text{user}}_{1}$): Appointment Schedule Secondary intent ($I^{\text{user}}_{2}$): Timing Rejection Appropriate assistant response: No problem. We have several other time slots available this week. Would Thursday morning at 9:30 AM or Friday afternoon at 3:15 PM work better for your schedule? Scenario 2. Refusing a service center location Assistant: Your vehicle’s compliance sticker needs replacement to avoid penalties. You can bring your vehicle to our main service center on Industrial Boulevard for the replacement. User: That doesn’t work for me. Primary intent ($I^{\text{user}}_{1}$): Appointment Schedule Secondary intent ($I^{\text{user}}_{2}$): Location Rejection Appropriate assistant response: I understand that location might be inconvenient. We also have service centers in Westview Mall and Northside Plaza.
@@
For each dialogue turn, we classify both assistant and user utterances using a two-stage intent classification process powered by Claude 3.7 Sonnet, shown in Algorithm 1.
Input: Dialogue turn $T$ with current utterances and dialogue history
foreach participant $P\in\{\text{assistant},\text{user}\}$ do
  Stage 1: Primary Intent Classification
  Construct prompt with:
  - Domain context
  - Dialogue history
  - Current dialogue turn
  - List of primary intents;
  Send prompt to Claude 3.7 Sonnet;
  Extract primary intent $I^{P}_{1}$ from response;
 1em Stage 2: Secondary Intent Classification
  Construct prompt with:
  - Domain context
  - Dialogue history
  - Current dialogue turn
  - Primary intent $I_{1}^{P}$ determined in Stage 1
  - Candidate secondary intents specific to $I_{1}^{P}$;
  Send prompt to Claude 3.7 Sonnet;
  Extract secondary intent $I^{P}_{2}$ from response;
end foreach
Algorithm 1 Two-Stage Intent Classification for Dialogue Turns
@@ B.1 Semantic Matching Algorithm
    $\begin{split}\text{Candidates}=&\{P_{i}\mid P_{i}.I^{\text{assistant}}_{1}==P.I^{\text{assistant}}_{1}\\ &\text{ AND }P_{i}.I^{\text{user}}_{1}==P.I^{\text{user}}_{1}\}\end{split}$ (5)
@@
    $\begin{split}P_{\text{selected}}=&\underset{i}{\text{argmax}}(\text{sim}(P,P_{i}))\\ &\text{ if }\max(\text{sim}(P,P_{i}))>\tau\end{split}$ (7)
@@ B.2 Effectiveness Evaluation
We use LLM-as-Judge to evaluate the generated response quality. Figure 3 shows that semantic matching significantly outperforms exact matching across all weight configurations, confirming the importance of this fallback mechanism in practical deployments.
@@ Appendix C Conversation-Based Retrieval Implementation Details
Appendix D Dataset Example
@@
Appendix E LLM-as-Judge Evaluation Prompt
For reproducibility, we provide the complete prompt template used for LLM-as-Judge evaluation:
Appendix F LLM-as-Judge Win Counts
To better illustrate the overall performance of different systems in our comparative evaluation, Figure 4 presents the win counts from LLM-as-Judge evaluations for both retrieval quality and response generation quality. A "win" is recorded when a system receives the highest rating among all compared systems for a particular test case. When multiple systems receive identical highest ratings, all are counted as winners.
The results clearly demonstrate CID-GraphRAG’s superior performance, with 73 wins in retrieval quality and 60 wins in response generation quality. For retrieval quality, CID-GraphRAG outperforms the second-best system (Conversation RAG) by 43%, showing the effectiveness of our dual-pathway approach in retrieving relevant context. For response generation, CID-GraphRAG achieves 58% more wins than the second-best system, demonstrating how improved retrieval translates to better response generation.
Figure 4: Performance comparison of different methods based on LLM-as-Judge win counts. The left group shows retrieval quality wins, and the right group shows response generation wins. CID-GraphRAG consistently outperforms all baseline methods in both aspects, with particularly significant advantages in retrieval quality.
Appendix G Computational Efficiency and Trade-offs
While CID-GraphRAG delivers substantial improvements in response quality, it introduces computational overhead compared to baseline approaches. This section analyzes these trade-offs and presents optimization strategies that make the approach viable for practical deployment.
Approach Token Latency (ms)
Intent Generation Intent Retrieval &
Recognition Recognition Generation
CID-GraphRAG 2136 1694 2649 2501
CID-GraphRAG (optimized) - 1694 11 2501
Intent RAG 2136 1542 2649 2467
Conversation RAG - 1702 - 2529
Direct LLM - 1203 - 2320
The primary computational cost in all evaluated approaches stems from LLM inference. CID-GraphRAG and Intent RAG requires additional LLM calls for intent recognition, resulting in significant token consumption and latency compared to semantic-only RAG baselines.
Thus, we propose an optimization strategy with lightweight intent recognition. Replacing LLM-based intent recognition with a fine-tuned lightweight model (e.g., BERT) could reduce intent recognition latency from 3 seconds to approximately 0.01 seconds, bringing CID-GraphRAG’s total latency to 2.5 seconds - comparable to direct LLM response generation while maintaining most quality advantages.

### v4

@@ 4.5 Limitations and Future Work
Domain Adaptability: Our current evaluation focuses on a specific domain of vehicle compliance-related customer service on a focused dataset of 268 conversations. Future work should systematically investigate the framework’s transferability to broader multi-turn dialogue scenarios, including complex task-oriented dialogues and cross-domain interactions where conversations span multiple service areas, i.e., well-recognized large-scale datasets such as MultiWOZ.
@@ 5 Conclusion
To address the challenges in multi-turn customer service dialogue systems, we propose CID-GraphRAG, a novel framework that integrates intent-driven graph structures with semantic similarity retrieval mechanisms. Our approach leverages knowledge graphs to enhance reasoning capabilities and enable more precise retrieval, while systematically incorporating dialogue history to improve contextual understanding and response coherence, thereby significantly boosting answer quality and relevance.
Experimental results demonstrate that CID-GraphRAG significantly outperforms both semantic-only and intent-only baseline approaches across all evaluation criteria, achieving a $57.9\%$ improvement in response quality over conventional RAG methods according to LLM-as-a-Judge head-to-head evaluations. These findings establish CID-GraphRAG as an effective framework for advancing conversational AI in knowledge-intensive domains where both contextual relevance and goal-oriented progression are essential. Future work will extend its applicability to broader multi-turn dialogue scenarios, explore the adaptability of intent hierarchy and integration of reinforcement learning.
Appendix A Intent Recognition Examples
@@
Scenario 1. Refusing a proposed appointment time Assistant: I see your vehicle compliance sticker expired last week. We have an opening tomorrow at 2:00 PM to replace it and bring your vehicle back into compliance. User: That doesn’t work for me. Primary intent ($I^{\text{user}}_{1}$): Appointment Schedule Secondary intent ($I^{\text{user}}_{2}$): Timing Rejection Appropriate assistant response: No problem. We have several other time slots available this week. Would Thursday morning at 9:30 AM or Friday afternoon at 3:15 PM work better for your schedule? Scenario 2. Refusing a service center location Assistant: Your vehicle’s compliance sticker needs replacement to avoid penalties. You can bring your vehicle to our main service center on XYZ Boulevard for the replacement. User: That doesn’t work for me. Primary intent ($I^{\text{user}}_{1}$): Appointment Schedule Secondary intent ($I^{\text{user}}_{2}$): Location Rejection Appropriate assistant response: I understand that location might be inconvenient. We also have service centers in A Mall and B Plaza.
@@ B.1 Semantic Matching Algorithm
    $\begin{split}\text{Candidates}=\{P_{i}\mid P_{i}.I^{\text{assistant}}_{1}==P.I^{\text{assistant}}_{1}\text{ AND }P_{i}.I^{\text{user}}_{1}==P.I^{\text{user}}_{1}\}\end{split}$ (5)
@@
    $\begin{split}P_{\text{selected}}=\underset{i}{\text{argmax}}(\text{sim}(P,P_{i}))\text{ if }\max(\text{sim}(P,P_{i}))>\tau\end{split}$ (7)
@@ B.2 Effectiveness Evaluation
We use LLM-as-a-Judge to evaluate the generated response quality. Figure 4 shows that semantic matching significantly outperforms exact matching across all weight configurations, confirming the importance of this fallback mechanism in practical deployments.
@@ Appendix C Conversation-Based Retrieval Implementation Details
Appendix D LLM-as-a-Judge Evaluation Prompt
For reproducibility, we provide the complete prompt template used for LLM-as-a-Judge evaluation:
Appendix E Dataset Example